OpenStudy (anonymous):

Last week Bill made a purchase of $56.78 before tax. This week, the same items on sale would have cost him $41.90 before tax. If the tax is 4%, how much could Bill have saved by buying the items on sale (including tax)?

OpenStudy (anonymous):

First you have to figure out the total for both items. This is solved by multiplying by .04 (4%) then adding the price. (56.78 * .04) + 56.78 2.27 + 56.78 = 59.05 (41.90 * .04) + 41.90 1.68 + 41.90 = 43.58 Now we have the totals you need to figure out how much he would of saved so you subtract the two totals. 59.05 - 43.58 = $15.47 This is how much he would have saved. my medal